Use only the slope and y-intercept to graph the following equation.
3x = 2y

-AND-

x + 5y = 5


I need to find out the slope and y-intercept of the two above questions, and how you got it?.....PLEASE....needed before 7:45 eastern time.

Answer by stanbon(75887)

Use only the slope and y-intercept to graph the following equation.
3x = 2y
y = (3/2)x + 0
slope = 3/2
y-intercept is 0
-----------------
Graph:
1st: plot the y-intercept which is (0,0)
2nd: From that point move your pencil one to the right and then (3/2) up
to the point (1,3/2)
3rd: Draw a line thru the two points.
----------------------------------------------
-AND-
x + 5y = 5
y = (-1/5)x + 1
slope is -1/5
y-intercept is 1
1st: plot the y-interept, whichis (0,1)
2nd From that point move your pencil one to the right and 1/5 downto
the point (1,4/5)
3rd: Draw a line thru the two points.
-------------------------------------------------
Cheers,
Stan H.